# Zilber’s Trichotomy in Hausdorff Geometric Structures

### v1

@@ 9.6. Smoothness and Enough Open Maps
We note, for example, that the hypothesis of Lemma 9.38 holds of the complex field with the analytic topology (though technically, this structure is not an éz topological field expansion, because it has no invariant basis). Indeed, in this case, the induced map $g:V(K)\rightarrow W(K)$ is a holomorphic map of $d$-dimensional complex manifolds with all fibers discrete, and one can apply a usual open mapping theorem from complex analysis (e.g. [GR84], p. 107). This was one of the key geometric properties used in [Cas23]), and is the inspiration for Lemma 9.38. Indeed, we will see later (Corollary 9.60) that the hypothesis of the lemma holds in ACVF in all characteristics.
@@ 10. Definable Slopes in ACVF
Otherwise, we use very little from the theory ACVF, so we do not elaborate here. Interested readers may refer to [HHM06, §2.1], [HK06, §2.1] and references therein for more background. Section 2.1 of the more recent [HOP24] contains also some background (and relevant references) on the analytic theory of functions definable in complete models of ACVF that is needed in a couple of points below.
@@ 10.1. Taylor Groupoids
    A $\emptyset$-ind-definable group $G$.
@@
Let $\mathcal{C}$ be a weak Taylor groupoid, and let $f:x\rightarrow y$ be a morphism in $\mathcal{C}$. For each $n\geq 1$, we define the $n$th truncation of $f$, dented $T_{n}(f)$, to be the tuple $(x,y,a(f),c_{1}(f),...,c_{n}(f))$.
@@ 10.5. TIMI in ACVF
We are given two $A$-definable families of correspondences in $K$, $X\subset K^{2}\times T$ and $Y\subset K^{2}\times U$, and generic points over $A$, $(x_{0},y_{0},t_{0})\in X$ and $(x_{0},y_{0},u_{0})\in Y$. We assume that the $n$-slopes of $X_{t_{0}}$ and $Y_{u_{0}}$ coincide at $(x_{0},y_{0})$, but that in some neighborhood of $u_{0}$, if $u\models\operatorname{tp}(u_{0}/x_{0}y_{0})$ with $u$ independent from $t_{0}$ over $x_{0}y_{0}$, then the $n$th slopes of $X_{t_{0}}$ and $Y_{u}$ do not coincide at $(x_{0},y_{0})$. We want to show that $(x_{0},y_{0},t_{0},u_{0})$ is a topological ramification point of the projection $I\rightarrow T\times U$, where $I$ is the set of $(x,y,t,u)$ with $(x,y)\in X_{t}\cap Y_{u}$.
@@
Now by the genericity of $(x_{0},y_{0},t_{0})\in X$ and $(x_{0},y_{0},u_{0})\in Y$, after restricting to small enough neighborhoods we can assume $X$ and $Y$ are defined by polynomial equations $P(x,y,t)=0$ and $P(x,y,u)=0$. Since the generalized étale functions of $X_{t_{0}}$ and $Y_{u_{0}}$ have no Frobenius power, it follows that the partial derivatives of $P$ in $Q$ in the $x$ and $y$ variables are non-zero at $(x_{0},y_{0},t_{0},u_{0})$. It follows that we are in the situation of Lemma 10.30: indeed, by compactness, any neighborhood of $u_{0}$ contains some $u\models\operatorname{tp}(u/x_{0}y_{0})$ with $u$ independent from $t_{0}$ over $x_{0}y_{0}$ – thus, by assumption, $(x_{0},y_{0},t_{0},u_{0})$ belongs to the closure of points $(x,y,t,u)\in I$ where the $n$th slopes of $X_{t}$ and $Y_{u}$ do not coincide. ∎
@@ 11. Proofs of the Main Theorems
We now prove our main theorems for definable relics of ACVF. For this, we need to recall the main result of the recent preprint [HOP24] (Fact 11.3 below). We use some ad hoc terminology to recall their setting most easily.
@@
The definition of local equivalence in [HOP24] is more general. It does not assume almost embeddability into $K$, and uses the more general $dp$-rank (discussed in the next section) instead of dimension. However, we will only encounter groups almost embeddable into $K$; and for these groups, the two notions coincide (see [Sim14, Theorem 0.3], and use [DGL11] to apply it).
@@
  • Case 1: $n_{0}=2$. Then $x+h_{13}x^{n_{0}}$ is the $n_{0}$-truncated composition of $x+h_{12}x^{n_{0}}$ and $x+h_{23}x^{n_{0}}$. One easily checks this truncated composition to be $x+(h_{12}+h_{23})^{n_{0}}$, and thus $h_{13}=h_{12}+h_{23}$.
@@
First, we check that we can assume $(K,v)$ is $\aleph_{1}$-saturated. This is a straightforward reduction, and similar arguments are given in [Cas23, Theorem 9.6] and [HS24, Lemma 4.8]. We give a sketch. Let $(K_{1},v_{1})$ be an $\aleph_{1}$-saturated elementary extension of $(K,v)$. Then the same formulas defining $\mathcal{M}$ in $(K,v)$ also define a definable $(K_{1},v_{1})$-relic, say $\mathcal{M}_{1}$. It is easy to see that $\mathcal{M}_{1}$ is an elementary extension of $\mathcal{M}$, so is strongly minimal and not locally modular.
@@
By elimination of imaginaries, we may assume $\mathcal{M}$ is a definable $K$-relic (that is, its universe is a subset of some $K^{n}$). By an identical argument to Theorem 11.5 above, we may assume $K$ is $\aleph_{1}$-saturated (really, $\aleph_{0}$ is enough here), and thus there is a non-trivial valuation on $K$. Let $v$ be such a valuation. Then $\mathcal{M}$ is still interpreted in $(K,v)$. One can now finish in two ways: either apply Theorem 11.5 directly (thus using [HOP24]); or apply Theorem 8.1 to show $\dim(M)=1$, and then conclude with the main result of [HS24]. ∎
@@ 12.3. Local Interalgebraicity
We will need the following stronger result (see [HOP24, Lemma 2.8]):
@@ 12.6. Corollaries of the Main Proposition
Fix $k$. We will find a $k$-rich configuration $(a,b,A)$ where $a\in X^{2}$ and $b$ is a tuple from $X$. First let $j=k+2$. Then let $\mathcal{C}=\{C_{t}:t\in T\}$ be an almost faithful family of plane curves in $\mathcal{M}$, where $T\subset M^{j+2}$ is generic (see [Cas23], Fact 2.27 and Lemma 2.42). Adding parameters, we assume $\mathcal{C}$ is $\mathcal{M}$-definable over $\emptyset$. Given $z\in M^{j}$, we let $\mathcal{C}(z)$ be the family with graph $\{(x,y,s)\in M\times M\times M^{2}:(x,y)\in C_{(s,z)}\}$. If $z$ is $\mathcal{M}$-generic in $M^{j}$, then $\mathcal{C}(z)$ is an almost faithful family indexed by a generic subset of $M^{2}$.
@@
Let $X\subset M$ be $D$-critical. Fix a rank 2 family $\mathcal{C}=\{C_{t}:t\in T\}$ of non-trivial plane curves as in Proposition 12.13, and assume it is $\emptyset$-definable in $\mathcal{M}$. Now let $v$ be any element of $M-\operatorname{acl}(0)$. Let $u\in X$ be such that $\operatorname{dprk}(u/v)=\operatorname{dprk}(X)$, and let $(x,y)\in X^{2}$ be such that $\operatorname{dprk}(xy/uv)=2\operatorname{dprk}(X)$. So $(u,v)$ is $\mathcal{M}$-generic in $M^{2}$. Then by Proposition 12.13, there is $t\in T$ with $(u,v),(x,y)\in C_{t}$ and $C_{t}\cap X^{2}$ infinite. Let $(x^{\prime},y^{\prime})\in C_{t}\cap X^{2}$ with $\operatorname{dprk}(x^{\prime}y^{\prime}/txy)\geq 1$. So $(x^{\prime},y^{\prime})$ is $\mathcal{M}$-generic in $C_{t}$ over $txy$, and thus there are only finitely many curves from $\mathcal{C}$ containing both $(x^{\prime},y^{\prime})$ and $(x,y)$ (see [Cas23], Lemma 2.38). Since $C_{t}$ is one such curve, $t\in\operatorname{acl}(xyx^{\prime}x^{\prime})$. But since $C_{t}$ is non-trivial, we also have $v\in\operatorname{acl}(tu)$. Thus in total we have $v\in\operatorname{acl}(uxyx^{\prime}y^{\prime})$. Now $u,x,y,x^{\prime},y^{\prime}\in X$. By compactness, all but finitely many elements of $M$ are in $\operatorname{acl}(X)$. Adding finitely many parameters to cover the remaining elements, we get that $M$ is almost internal to $X$. But by definition $X$ is almost internal to $D$, so it follows that $M$ is almost internal to $D$. ∎

### v2

@@ 9.6. Smoothness and Enough Open Maps
We note, for example, that the hypothesis of Lemma 9.38 holds of the complex field with the analytic topology (though technically, this structure is not an éz topological field expansion, because it has no invariant basis). Indeed, in this case, the induced map $g:V(K)\rightarrow W(K)$ is a holomorphic map of $d$-dimensional complex manifolds with all fibers discrete, and one can apply a usual open mapping theorem from complex analysis (e.g. [GR84], p. 107). This was one of the key geometric properties used in [Cas24b]), and is the inspiration for Lemma 9.38. Indeed, we will see later (Corollary 9.60) that the hypothesis of the lemma holds in ACVF in all characteristics.
@@ 10. Definable Slopes in ACVF
Otherwise, we use very little from the theory ACVF, so we do not elaborate here. Interested readers may refer to [HHM06, §2.1], [HK06, §2.1] and references therein for more background. Section 2.1 of the more recent [HOP25] contains also some background (and relevant references) on the analytic theory of functions definable in complete models of ACVF that is needed in a couple of points below.
@@ 10.1. Taylor Groupoids
    A $\emptyset$-ind-definable group $G$ contained in $\operatorname{dcl}(\emptyset)$.
@@
Let $\mathcal{C}$ be a weak Taylor groupoid, and let $f:x\rightarrow y$ be a morphism in $\mathcal{C}$. For each $n\geq 0$, we define the $n$th truncation of $f$, dented $T_{n}(f)$, to be the tuple $(x,y,a(f),c_{1}(f),...,c_{n}(f))$ (if $n=0$ this just means $(x,y,a(f))$).
@@ 10.5. TIMI in ACVF
We are given two $A$-definable families of correspondences in $K$, $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ (with graphs $X\subset K^{2}\times T$ and $Y\subset K^{2}\times U$), and a generic $(\mathcal{X},\mathcal{Y})$-tangency $(x_{0},y_{0},t_{0},u_{0},n)$ over $A$. We want to show that $(x_{0},y_{0},t_{0},u_{0})$ is a topological ramification point of the projection $I\rightarrow T\times U$, where $I$ is the set of $(x,y,t,u)$ with $(x,y)\in X_{t}\cap Y_{u}$.
@@
Now by the genericity of $(x_{0},y_{0},t_{0})\in X$ and $(x_{0},y_{0},u_{0})\in Y$, after restricting to small enough neighborhoods we can assume $X$ and $Y$ are defined by polynomial equations $P(x,y,t)=0$ and $P(x,y,u)=0$. Since the generalized étale functions of $X_{t_{0}}$ and $Y_{u_{0}}$ have no Frobenius power, it follows that the partial derivatives of $P$ in $Q$ in the $x$ and $y$ variables are non-zero at $(x_{0},y_{0},t_{0},u_{0})$. It follows that we are in the situation of Lemma 10.30: indeed, by assumption $(x_{0},y_{0},t_{0},u_{0})$ is strongly approximable over $A$ and $\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching over $A$ in a neighborhood of $(x_{0},y_{0},t_{0},u_{0})$ – thus $(x_{0},y_{0},t_{0},u_{0})$ belongs to the closure of points $(x,y,t,u)\in I$ (namely strongly generic $(\mathcal{X},\mathcal{Y})$-intersections) where the $n$th slopes of $X_{t}$ and $Y_{u}$ do not coincide. ∎
@@ 11. Proofs of the Main Theorems
We now prove our main theorems for definable relics of ACVF. For this, we need to recall the main result of the recent preprint [HOP25] (Fact 11.3 below). We use some ad hoc terminology to recall their setting most easily.
@@
The definition of local equivalence in [HOP25] is more general. It does not assume almost embeddability into $K$, and uses the more general $dp$-rank (discussed in the next section) instead of dimension. However, we will only encounter groups almost embeddable into $K$; and for these groups, the two notions coincide (see [Sim14, Theorem 0.3], and use [DGL11] to apply it).
@@
  • Case 2: $n_{0}>1$. Then $x+h_{13}x^{n_{0}}$ is the $n_{0}$-truncated composition of $x+h_{12}x^{n_{0}}$ and $x+h_{23}x^{n_{0}}$. One easily checks this truncated composition to be $x+(h_{12}+h_{23})^{n_{0}}$, and thus $h_{13}=h_{12}+h_{23}$.
@@
First, we check that we can assume $(K,v)$ is $\aleph_{1}$-saturated. This is a straightforward reduction, and similar arguments are given in [Cas24b, Theorem 9.6] and [HS25, Lemma 4.8]. We give a sketch. Let $(K_{1},v_{1})$ be an $\aleph_{1}$-saturated elementary extension of $(K,v)$. Then the same formulas defining $\mathcal{M}$ in $(K,v)$ also define a definable $(K_{1},v_{1})$-relic, say $\mathcal{M}_{1}$. It is easy to see that $\mathcal{M}_{1}$ is an elementary extension of $\mathcal{M}$, so is strongly minimal and not locally modular.
@@
By elimination of imaginaries, we may assume $\mathcal{M}$ is a definable $K$-relic (that is, its universe is a subset of some $K^{n}$). By an identical argument to Theorem 11.5 above, we may assume $K$ is $\aleph_{1}$-saturated (really, $\aleph_{0}$ is enough here), and thus there is a non-trivial valuation on $K$. Let $v$ be such a valuation. Then $\mathcal{M}$ is still interpreted in $(K,v)$. One can now finish in two ways: either apply Theorem 11.5 directly (thus using [HOP25]); or apply Theorem 8.1 to show $\dim(M)=1$, and then conclude with the main result of [HS25]. ∎
@@ 12.3. Local Interalgebraicity
We will need the following stronger result (see [HOP25, Lemma 2.8]):
@@ 12.6. Corollaries of the Main Proposition
Fix $k$. We will find a $k$-rich configuration $(a,b,A)$ where $a\in X^{2}$ and $b$ is a tuple from $X$. First let $j=k+2$. Then let $\mathcal{C}=\{C_{t}:t\in T\}$ be an almost faithful family of plane curves in $\mathcal{M}$, where $T\subset M^{j+2}$ is generic (see [Cas24b], Fact 2.27 and Lemma 2.42). Adding parameters, we assume $\mathcal{C}$ is $\mathcal{M}$-definable over $\emptyset$. Given $z\in M^{j}$, we let $\mathcal{C}(z)$ be the family with graph $\{(x,y,s)\in M\times M\times M^{2}:(x,y)\in C_{(s,z)}\}$. If $z$ is $\mathcal{M}$-generic in $M^{j}$, then $\mathcal{C}(z)$ is an almost faithful family indexed by a generic subset of $M^{2}$.
@@
Let $X\subset M$ be $D$-critical. Fix a rank 2 family $\mathcal{C}=\{C_{t}:t\in T\}$ of non-trivial plane curves as in Proposition 12.13, and assume it is $\emptyset$-definable in $\mathcal{M}$. Now let $v$ be any element of $M-\operatorname{acl}(0)$. Let $u\in X$ be such that $\operatorname{dprk}(u/v)=\operatorname{dprk}(X)$, and let $(x,y)\in X^{2}$ be such that $\operatorname{dprk}(xy/uv)=2\operatorname{dprk}(X)$. So $(u,v)$ is $\mathcal{M}$-generic in $M^{2}$. Then by Proposition 12.13, there is $t\in T$ with $(u,v),(x,y)\in C_{t}$ and $C_{t}\cap X^{2}$ infinite. Let $(x^{\prime},y^{\prime})\in C_{t}\cap X^{2}$ with $\operatorname{dprk}(x^{\prime}y^{\prime}/txy)\geq 1$. So $(x^{\prime},y^{\prime})$ is $\mathcal{M}$-generic in $C_{t}$ over $txy$, and thus there are only finitely many curves from $\mathcal{C}$ containing both $(x^{\prime},y^{\prime})$ and $(x,y)$ (see [Cas24b], Lemma 2.38). Since $C_{t}$ is one such curve, $t\in\operatorname{acl}(xyx^{\prime}x^{\prime})$. But since $C_{t}$ is non-trivial, we also have $v\in\operatorname{acl}(tu)$. Thus in total we have $v\in\operatorname{acl}(uxyx^{\prime}y^{\prime})$. Now $u,x,y,x^{\prime},y^{\prime}\in X$. By compactness, all but finitely many elements of $M$ are in $\operatorname{acl}(X)$. Adding finitely many parameters to cover the remaining elements, we get that $M$ is almost internal to $X$. But by definition $X$ is almost internal to $D$, so it follows that $M$ is almost internal to $D$. ∎